OpenStudy (anonymous):

At room temperature, the average speed of an oxygen molecule is about 500 m/s. What is this in miles per hour? (1 kilometer = 0.62 miles)

OpenStudy (hwyl):

\(\large 500 \frac{m}{sec} \times \frac{1~km}{1000~m} \times \frac{0.62~mi}{1~km} \times \frac{60 ~sec}{1 ~min} \times \frac{60~min}{1 ~hr} \)

OpenStudy (anonymous):

1116, but the answer choices are a.10 miles/hr b.100 miles/hr c.1000 miles/hr d.10,000 miles/hr

OpenStudy (hwyl):

okay, and?

OpenStudy (anonymous):

so would that equal to 1000 mi/hr

OpenStudy (hwyl):

you tell me I already laid out the solution to you

OpenStudy (hwyl):

what i did is considered helping you cheat in fact

OpenStudy (anonymous):

its from a practice test that my teacher gave me. i already have the answer but idk why its 1000

OpenStudy (anonymous):

you dont need to accuse me for cheating.

OpenStudy (hwyl):

i already showed you how to solve for it